# SlideAVSR: A Dataset of Paper Explanation Videos for Audio-Visual Speech Recognition

### v1

@@ Abstract
[…]
@@ 2 Related Work
Compared to the efforts that have been made on lip reading datasets (Chung et al., 2017; Chung and Zisserman, 2017a, b; Afouras et al., 2018a, b; Shillingford et al., 2019), AVSR datasets in other types of videos remain scarce. To our knowledge, VisSpeech (Gabeur et al., 2022) and the audiovisual diarization benchmark in the Ego4D challenge (Jain et al., 2023) are the only AVSR datasets not centered around lip reading. VisSpeech is constructed from a subset of the instructional video dataset HowTo100M (Miech et al., 2019), where the visual stream and speech audio are semantically related. The audiovisual diarization benchmark in the Ego4D challenge consists of 585 egocentric video clips. It is imperative to build more diverse benchmark datasets to evaluate the image comprehension capabilities of AVSR models.
@@ 3.2 Filtering
We perform three times of generation, and if “Yes” is outputted at least once, we adopt the video; otherwise, we discard it. We show the details of the model and prompt in Appendix B.
@@
We perform generation for each screenshot, and if “Yes” is outputted at least once, we adopt the video; otherwise, we discard it. We show the details of the model and prompt in Appendix B.
@@ 3.3 Cleansing
Due to the inaccuracy in the timing of subtitles, we implement audio-subtitle alignment and scoring using CTC segmentation Kürzinger et al. (2020). We set the threshold to -7 and exclude utterances with lower scores. The details of the model are shown in Appendix B.
@@
We capture screenshots at the midpoint of each utterance, followed by filtering using BLIP-2. Three generations are conducted for each screenshot, and if “Yes” is outputted at least once, we adopt the utterance; otherwise, we discard it. The employed prompt is identical to the BLIP-2 filter in Section 3.2.
@@ 3.4 Data Partitioning
Previous studies (Meyer et al., 2020; Javed et al., 2023; DiChristofano et al., 2023) have suggested that the performance of ASR systems significantly varies depending on the speaker’s accent [^3] The term “accent” in this paper refers to comprehensive prosodic information, including accent, intonation, tone, etc.. Based on the hypothesis that visual information contributes to the recognition of challenging accents, we ask native English speakers to classify the speakers’ accents in SlideAVSR and perform dataset partitioning. We partition the dataset into Train, Dev, and TestA, reserving a smaller yet significant TestB subset for South Asian English (SAE) accents. During partitioning, we have ensured that the same speaker did not belong to multiple partitions. Additionally, videos with machine-generated audio were manually excluded by the annotators.
@@ 4.1 Approaches
4.2 Implement Details
We used Whisper large-v3555https://huggingface.co/openai/whisper-large-v3 as a base model and Word Error Rate (WER) for evaluation. In the case of DocWhisper, we captured screenshots at the midpoint of each utterance, fed them into the OCR module, and used the recognized text as the prompts to Whisper. We use Google Cloud Vision API [^6] https://cloud.google.com/vision for OCR. The prompts were presented to the model as word sequences, such as “word 1, word 2, …, word $n$”. FQ Ranker utilized word frequency counts obtained from the English Wikipedia as of April 2023 and sorted the OCR results in ascending order based on word frequency. We conducted experiments with different maximum word counts for prompts ($K\in\{25,50,75,100\}$) and with or without FQ Ranker. More implementation details are provided in Appendix C.
@@ 4.2 Implement Details
Model Modality Fine-tune $K$${}^{a}$ TestA TestB
@@
${}^{a}$ndicating maximum word counts for prompts.
@@ 5 Conclusion and Future Work
In comparison to mainstream AVSR datasets, SlideAVSR exhibits a notably limited number of speakers. This may lead to data imbalance and create obstacles to the model’s training process. In addition, due to our focused collection of scientific paper explanation videos related to artificial intelligence, imbalances may have emerged in terms of speaker nationality, age, and gender.
@@ Appendix B Models and prompts used in data filtering and cleansing
We used kamo-naoyuki_wsj [^8] https://huggingface.co/espnet/kamo-naoyuki_wsj and ESPnet implemenations [^9] https://github.com/espnet/espnet.

### v2

@@ Abstract
Hao Wang1  Shuhei Kurita2  Shuichiro Shimizu3  Daisuke Kawahara1,2,4 1 Waseda University  2 National Institute of Informatics (NII)  3 Kyoto University  4 LLMC, NII conan1024hao@akane.waseda.jpskurita@nii.ac.jp sshimizu@nlp.ist.i.kyoto-u.ac.jpdkw@waseda.jp
@@ 2 Related Work
Compared to the efforts that have been made on lip reading datasets (Chung et al., 2017; Chung and Zisserman, 2017a, b; Afouras et al., 2018a, b; Shillingford et al., 2019), AVSR datasets in other types of videos remain scarce. VisSpeech (Gabeur et al., 2022) is constructed from a subset of the instructional video dataset HowTo100M (Miech et al., 2019) where the visual stream and speech audio are semantically related. The audiovisual diarization benchmark in the Ego4D challenge (Jain et al., 2023) consists of 585 egocentric video clips. In terms of utilizing textual information extracted from videos, SlideSpeech Wang et al. (2023) builds an AVSR dataset on online conference videos enriched with slides. However, the filtering process of SlideSpeech relies heavily on human annotators. In this work, we exclude videos and utterances that do not correspond to slides by utilizing a vision language model. This approach reduces annotation costs and enhances the purity and quality of our dataset within the slide domain.
@@ 3.2 Filtering
We perform three times of generation, and if “Yes” is outputted at least once, we adopt the video; otherwise, we discard it. We show the details of the model and prompt in Appendix B. A total of 342 videos were excluded, leaving 294 videos for subsequent processes.
@@
We perform generation for each screenshot, and if “Yes” is outputted at least once, we adopt the video; otherwise, we discard it. We show the details of the model and prompt in Appendix B. A total of 6 videos were excluded, leaving 288 videos for subsequent processes.
@@
A total of 38 videos were excluded, leaving 250 videos for subsequent processes.
@@ 3.3 Cleansing
Due to the inaccuracy in the timing of subtitles, we implement audio-subtitle alignment and scoring using CTC segmentation Kürzinger et al. (2020). We set the threshold to -7 and exclude utterances with lower scores. The details of the model are shown in Appendix B. Approximately 2.5% of utterances were excluded through this process.
@@
We capture screenshots at the midpoint of each utterance, followed by filtering using BLIP-2. Three generations are conducted for each screenshot, and if “Yes” is outputted at least once, we adopt the utterance; otherwise, we discard it. The employed prompt is identical to the BLIP-2 filter in Section 3.2. Approximately 1.0% of utterances were excluded through this process.
@@ 3.4 Data Partitioning
Previous studies (Meyer et al., 2020; Javed et al., 2023; DiChristofano et al., 2023) have suggested that the performance of ASR systems significantly varies depending on the speaker’s accent [^3] The term “accent” in this paper refers to comprehensive prosodic information, including accent, intonation, tone, etc.. Based on the hypothesis that visual information contributes to the recognition of challenging accents, we ask native English speakers to classify the speakers’ accents in SlideAVSR and perform dataset partitioning. We partition the dataset into Train, Dev, and TestA, reserving a smaller yet significant TestB subset for South Asian English (SAE) accents. During partitioning, we have ensured that the same speaker did not belong to multiple partitions. Additionally, 5 videos with machine-generated audio were manually excluded by the annotators.
@@ 4.1 Approaches
Model Modality Fine-tune $K$a TestA TestB
@@
aIndicating maximum word counts for prompts.
@@
4.2 Implement Details
We used Whisper large-v3555https://huggingface.co/openai/whisper-large-v3 as a base model and Word Error Rate (WER) for evaluation. In the case of DocWhisper, we captured screenshots at the midpoint of each utterance, fed them into the OCR module, and used the recognized text as the prompts to Whisper. In this case, multiple utterances might share the same slide. We use Google Cloud Vision API [^6] https://cloud.google.com/vision for OCR. The prompts were presented to the model as word sequences, such as “word 1, word 2, …, word $n$”. FQ Ranker utilized word frequency counts obtained from the English Wikipedia and sorted the OCR results in ascending order based on word frequency. We conducted experiments with different maximum word counts for prompts ($K\in\{25,50,75,100\}$) and with or without FQ Ranker. More implementation details are provided in Appendix C.
@@ 5 Conclusion and Future Work
In comparison to mainstream AVSR datasets, SlideAVSR exhibits a notably limited number of videos and speakers. This may lead to data imbalance and create obstacles to the model’s training process. In addition, due to our focused collection of scientific paper explanation videos related to artificial intelligence, imbalances may have emerged in terms of speaker nationality, age, and gender.
Compared to SlideSpeech Wang et al. (2023), we introduced a vision-language model to filter videos and utterances that do not correspond to the slides, thereby reducing annotation costs and improving the quality of the dataset. However, our dataset construction process still relies on manual annotation. Fully automating this process will be a major challenge for the future.
@@ Appendix B Models and prompts used in data filtering and cleansing
We used kamo-naoyuki_wsj [^8] https://huggingface.co/espnet/kamo-naoyuki_wsj and ESPnet implemenations [^9] https://github.com/espnet/espnet